# The ambiguous AT2022rze: Changing-look AGN mimicking a supernova in a merging galaxy system

### v1

@@ 4.1 SLSNe II
AT2022rze was originally classified as a H-rich SN based on the presence of narrow Balmer lines. We show in Sect. 3.3 that the position of the transient (galaxy B, see Fig. 1) may be associated to an AGN. This association does not exclude the possibility of a SN. On the contrary, it has been proposed that the rates of CCSNe are the same or even higher in AGNs or star-forming galaxies (Petrosian & Turatto, 1990; Petrosian et al., 2005).
In SNe, the presence of narrow lines indicate CSM interaction. Its bumpy light curve is also consistent with the presence of CSM, and it is reminiscent of other long-lived H-rich events (see Sect. 3.4). It has been argued that these bumpy, long-lived light curves result from the interaction between the SN ejecta and a CSM with large density variations. Such a CSM structure can be produced by episodic mass-loss events. These are often associated with Luminous Blue Variable-like progenitors (e.g., Smith et al., 2011) or Pulsational Pair Instability SNe (PPISNe, e.g: Woosley, 2017, 2018).
At early times, the observed colors of AT2022rze agree more with those observed for SLSNe II than for TDEs (see Fig. 4). The estimated total radiated energy (see Sect. 3.1.3) can be explained by CSM interaction. However, a rough analysis of the impact of the potential host extinction indicates that the transient may be rather blue (see Sect. 3.3) and much more energetic, reaching energies of the order of erg. This would present a challenge to a supernova interpretation, as the radiated energy would exceed what is available in a neutrino-driven core-collapse explosion (Janka, 2012). In this case, the rotational energy of a central engine would be required, in addition to CSM interaction to produce the spectral signatures and the structure in the light curve (Woosley, 2017).
The spectra of AT2022rze show several coronal lines (see Sect. 3.2), with [Fe vii] 6087, [Fe x] 6374, and [Fe xi] 7892 being the strongest ones. Coronal lines are uncommon in SN although not previously unseen in SN IIn (e.g., Fransson et al., 2002; Smith et al., 2009; Izotov & Thuan, 2009), where X-rays from the SN excite the CSM. Yet, it has been argued that coronal line emitting transients are associated to TDEs (e.g., Hinkle et al., 2024) and AGN (e.g., Rose et al., 2015).
@@ 4.2 TDE
TDEs are characterized for their blue colors and broad Balmer lines without a blueshifted absorption. AT2022rze show red colors unless we consider high host extinction (see Fig.4). In addition, The H line are much narrower than those in other TDEs (see Fig. 12). We calculate the time to and from half-peak (t and t, respectively), and compare them to the same values presented for TDEs by Yao et al. (2023). We find that the half-peak rise and decline time of AT2022rze is similar to those of TDEs, with AT2022rze being among the TDEs with the longest rest-frame duration above half-peak. We note, however, that it is possible that galaxy B hosts an AGN, and it is unclear what the characteristics of the TDE associated to an AGN should look like (Zabludoff et al., 2021).
The host galaxy of AT2022rze is experiencing a merger. Although a few TDEs have been suggested to be associated to merging galaxies (e.g., Mattila et al., 2018; Kool et al., 2020; Reynolds et al., 2022), there are only two confirmed TDEs in optical wavelenghts in such systems (Charalampopoulos et al., 2024; Onori et al., 2025). These events exhibit much smoother light curves and spectroscopic features that differ significantly from those of AT2022rze, making direct comparison unwarranted.
@@
It is often difficult to differentiate between CLAGNs and TDEs due to their overlapping observational features. It has even been proposed that CLAGN are triggered by TDEs (Merloni et al., 2015; Wang et al., 2024). The analysis of the DESI spectrum of galaxy B, together with the Blast analysis, suggest that galaxy B is associated with an AGN.
@@
Using C-statistics (Cash, 1979) to model the eROSITA spectrum with an absorbed power-law (tbabs*cflux*powerlaw in xspec, see Fig. 13), with the Galactic column density fixed at (HI4PI Collaboration et al., 2016), we find a decent fit with cstat/dof=18.1/11, a photon index of and an unabsorbed 0.3–2.3 keV flux of . The latter corresponds to an X-ray luminosity of , which is on the fainter end of AGN, consistent with low-luminosity or obscured AGN (Hasinger et al., 2005).
@@
There are several sub-classifications of AGN, in particular, they can be divided in radio-loud and radio-quiet (Kellermann et al., 1989). VLASS archival pre-transient images put sub-mJy upper limits near the center of the host (see Sect. 2), but there is no radio detection of galaxy B. This suggests a radio-quiet AGN (e.g: Padovani et al., 2009). Radio-quiet AGN are often associated to spiral galaxies, and share many characteristics with star forming galaxies (Bonzini et al., 2013).
@@
Given the evidence of the presence of an AGN at the center of galaxy B, it is possible that AT2022rze is associated to changes on the environment or accretion rate of said AGN. Considering the transition from the AGN (galaxy B) to the LINER region (earlier spectra of AT2022rze, see Fig. 11), AT2022rze could be considered a Changing-look AGN (CLAGN). CLAGN is a fairly recent and slowly growing class of transients, but the number of events is still low. Their characteristics are very diverse, and their searches can be considered as an anomaly detection task (e.g., Sánchez-Sáez et al., 2021). Therefore, systematic analysis of their time-scales and overall light curve and spectral characteristics is scarce. López-Navas et al. (2023) developed a method to select CLAGN using multi-wavelength observations. Unfortunately, we do not have multi-wavelength coverage of AT2022rze itself. Still, the host X-ray flux measured before the transient went off, is consistent with the pre-activity X-ray fluxes that López-Navas et al. (2023) measure for their sample of CLAGNs. In addition, the evolution of W1W2 color at the position of AT2022rze that goes from < 0.5 mag before it is detected (see Table 1) to 1.1 mag (see last column of Table 4) for the last observed epoch, is in agreement with the interpretation presented by López-Navas et al. (2023), who argue that this effect is produced by a stronger contribution of hot dust produced by the increase in brightness of the AGN.
@@
We argue in Sect. 3.3 that the steep Balmer decrement values are associated to significant host extinction. However, there is evidence that the Balmer decrement can depend on the accretion-rate for some AGNs and CLAGNs, in such cases the steep Balmer decrement could be related to ionizing photons with low flux (Wu et al., 2023). The fact that the Balmer decrement evolves with time (see Fig. 12), suggests that other mechanisms are in place in addition to extinction. Still, the fact that coronal lines can be seen during the whole spectra evolution, suggest the presence of high energies being in place.
@@ 5 Conclusions
We have studied AT2022rze to try to determine its nature. The host of AT2022rze seems to be formed of two galaxies (A and B) undergoing a merging event. The transient is located in the center of galaxy B, located South-East of the plane-of-sky geometrical center of the system. AT2022rze exhibits minimal spectral evolution, consistently displaying narrow Balmer lines alongside various high-excitation features, indicative of a persistent high-energy source. The X-ray luminosity observed near the position of the transient before it went off is much lower than that observed in AGNs (e.g: Elvis et al., 1978, 1990; Serlemitsos et al., 1996), although we note that the observations are off-set from the position of the transient. Coronal lines present in the spectra of AT2022rze suggest the presence of soft X-ray and UV emission. As coronal lines are found in AGNs, TDEs, and SNe, their presence alone do not serve as a diagnostic. The overall observed spectroscopic and light curve features of the transient are reminiscent of SLSNe II, TDEs and AGNs.
@@
It is worth noting that the sustained presence of coronal lines in the spectra of some transients has led to the proposal of a new class, referred to as coronal line emitters (ECLs). Although this class groups together events that appear to exhibit similar phenomenology, it is not defined by a shared underlying physical mechanism, as the SLSN, TDE, or AGN classes are. Although ECLs have been mostly associated with TDEs (Hinkle et al., 2024), it is still not clear what is their association to CLAGNs (Clark et al., 2024, 2025).
We find possible evidence of Bowen resonance-fluorescence (see Sect. 3.2). The number of events displaying Bowen fluorescence features has been steadily increasing and a new class of transients dubbed Bowen fluorescence flares has been proposed (BFFs Trakhtenbrot et al., 2019). Again, this class is defined by observational characteristics rather than by an underlying physical mechanism, and the nature of these transients is still debated (e.g., Śniegowska et al., 2025).
AT2022rze is a luminous ambiguous transient in the nucleus of a merging galaxy. Many such ambiguous events have recently been found and have been dubbed ambiguous nuclear transients (ANTs, Wiseman et al., 2025), which is another phenomenological class. Wiseman et al. (2025) propose that ANTs could be TDEs produced by different mechanisms than that of standard TDEs.
@@ Acknowledgements
P.J.P thanks Stephen Thorp, Laureano Martinez, and Alberto Saldana-Lopez for useful discussion. Funded by the European Union (ERC, project number 101042299, TransPIre). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. CPG acknowledges financial support from the Secretary of Universities and Research (Government of Catalonia) and by the Horizon 2020 Research and Innovation Programme of the European Union under the Marie Skłodowska-Curie and the Beatriu de Pinós 2021 BP 00168 programme, from the Spanish Ministerio de Ciencia e Innovación (MCIN) and the Agencia Estatal de Investigación (AEI) 10.13039/501100011033 under the PID2023-151307NB-I00 SNNEXT project, from Centro Superior de Investigaciones Científicas (CSIC) under the PIE project 20215AT016 and the program Unidad de Excelencia María de Maeztu CEX2020-001058-M, and from the Departament de Recerca i Universitats de la Generalitat de Catalunya through the 2021-SGR-01270 grant. SM acknowledges support from the Research Council of Finland project 350458. FP acknowledges support from the Spanish Ministerio de Ciencia, Innovación y Universidades (MICINN) under grant numbers PID2022-141915NB-C21. This work is based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grants No. AST-1440341 and AST-2034437 and a collaboration including current partners Caltech, IPAC, the Weizmann Institute of Science, the Oskar Klein Center at Stockholm University, the University of Maryland, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, IN2P3, University of Warwick, Ruhr University Bochum, Northwestern University and former partners the University of Washington, Los Alamos National Laboratories, and Lawrence Berkeley National Laboratories. Operations are conducted by COO, IPAC, and UW. The ZTF forced-photometry service was funded under the Heising-Simons Foundation grant No. 12540303 (PI: Graham). SED Machine is based upon work supported by the National Science Foundation under Grant No. 1106171. The Gordon and Betty Moore Foundation, through both the Data-Driven Investigator Program and a dedicated grant, provided critical funding for SkyPortal. This work has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. The Asteroid Terrestrial-impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile. We acknowledge ESA Gaia, DPAC and the Photometric Science Alerts Team (http://gsaweb.ast.cam.ac.uk/alerts). This research has made use of the NASA/IPAC Extragalactic Database, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research has made use of the SVO Filter Profile Service "Carlos Rodrigo", funded by MCIN/AEI/10.13039/501100011033/ through grant PID2020-112949GB-I00 This research has made use of the SVO Filter Profile Service “Carlos Rodrigo”, funded by MCIN/AEI/10.13039/501100011033/ through grant PID2020-112949GB-I00 The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This research has made use of the CIRADA cutout service at cutouts.cirada.ca, operated by the Canadian Initiative for Radio Astronomy Data Analysis (CIRADA). CIRADA is funded by a grant from the Canada Foundation for Innovation 2017 Innovation Fund (Project 35999), as well as by the Provinces of Ontario, British Columbia, Alberta, Manitoba and Quebec, in collaboration with the National Research Council of Canada, the US National Radio Astronomy Observatory and Australia’s Commonwealth Scientific and Industrial Research Organisation. The Legacy Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS; Proposal ID 2014B-0404; PIs: David Schlegel and Arjun Dey), the Beijing-Arizona Sky Survey (BASS; NOAO Prop. ID 2015A-0801; PIs: Zhou Xu and Xiaohui Fan), and the Mayall z-band Legacy Survey (MzLS; Prop. ID 2016A-0453; PI: Arjun Dey). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory (LBNL). The Legacy Surveys project is honored to be permitted to conduct astronomical research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation.

### v2

@@ 4.1 SLSNe II
AT2022rze was originally classified as a H-rich SN based on the presence of narrow Balmer lines. We show in Sect. 3.3 that the position of the transient (galaxy B, see Fig. 1) may be associated to an AGN. This association does not exclude the possibility of a SN. On the contrary, it has been proposed that the rates of CCSNe are the same or even higher in AGNs or star-forming galaxies (1990A&A...239...63P; 2005AJ....129.1369P).
In SNe, the presence of narrow lines indicate CSM interaction. Its bumpy light curve is also consistent with the presence of CSM, and it is reminiscent of other long-lived H-rich events (see Sect. 3.4). It has been argued that these bumpy, long-lived light curves result from the interaction between the SN ejecta and a CSM with large density variations. Such a CSM structure can be produced by episodic mass-loss events. These are often associated with Luminous Blue Variable-like progenitors (e.g., 2011MNRAS.415..773S) or Pulsational Pair Instability SNe (PPISNe, e.g: 2017ApJ...836..244W; 2018ApJ...863..105W).
At early times, the observed colors of AT2022rze agree more with those observed for SLSNe II than for TDEs (see Fig. 4). The estimated total radiated energy (see Sect. 3.1.3) can be explained by CSM interaction. However, a rough analysis of the impact of the potential host extinction indicates that the transient may be rather blue (see Sect. 3.3) and much more energetic, reaching energies of the order of erg. This would present a challenge to a supernova interpretation, as the radiated energy would exceed what is available in a neutrino-driven core-collapse explosion (2012ARNPS..62..407J). In this case, the rotational energy of a central engine would be required, in addition to CSM interaction to produce the spectral signatures and the structure in the light curve (2017ApJ...836..244W).
The spectra of AT2022rze show several coronal lines (see Sect. 3.2), with [Fe vii] 6087, [Fe x] 6374, and [Fe xi] 7892 being the strongest ones. Coronal lines are uncommon in SN although not previously unseen in SN IIn (e.g., 2002ApJ...572..350F; 2009ApJ...695.1334S; 2009ApJ...707.1560I), where X-rays from the SN excite the CSM. Yet, it has been argued that coronal line emitting transients are associated to TDEs (e.g., 2024MNRAS.528.4775H) and AGN (e.g., 2015MNRAS.448.2900R).
@@ 4.2 TDE
TDEs are characterized for their blue colors and broad Balmer lines without a blueshifted absorption. AT2022rze show red colors unless we consider high host extinction (see Fig.4). In addition, The H line are much narrower than those in other TDEs (see Fig. 12). We calculate the time to and from half-peak (t and t, respectively), and compare them to the same values presented for TDEs by 2023ApJ...955L...6Y. We find that the half-peak rise and decline time of AT2022rze is similar to those of TDEs, with AT2022rze being among the TDEs with the longest rest-frame duration above half-peak. We note, however, that it is possible that galaxy B hosts an AGN, and it is unclear what the characteristics of the TDE associated to an AGN should look like (2021SSRv..217...54Z).
The host galaxy of AT2022rze is experiencing a merger. Although a few TDEs have been suggested to be associated to merging galaxies (e.g., 2018Sci...361..482M; 2020MNRAS.498.2167K; 2022A&A...664A.158R), there are only two confirmed TDEs in optical wavelenghts in such systems (2024A&A...689A.350C; 2025MNRAS.540..498O). These events exhibit much smoother light curves and spectroscopic features that differ significantly from those of AT2022rze, making direct comparison unwarranted.
@@
It is often difficult to differentiate between CLAGNs and TDEs due to their overlapping observational features. It has even been proposed that CLAGN are triggered by TDEs (2015MNRAS.452...69M; 2024arXiv240612096W). The analysis of the DESI spectrum of galaxy B, together with the Blast analysis, suggest that galaxy B is associated with an AGN.
@@
Using C-statistics (Cash1979) to model the eROSITA spectrum with an absorbed power-law (tbabs*cflux*powerlaw in xspec, see Fig. 13), with the Galactic column density fixed at (HI4PI2016), we find a decent fit with cstat/dof=18.1/11, a photon index of and an unabsorbed 0.3–2.3 keV flux of . The latter corresponds to an X-ray luminosity of , which is on the fainter end of AGN, consistent with low-luminosity or obscured AGN (Hasinger2005).
@@
There are several sub-classifications of AGN, in particular, they can be divided in radio-loud and radio-quiet (1989AJ.....98.1195K). VLASS archival pre-transient images put sub-mJy upper limits near the center of the host (see Sect. 2), but there is no radio detection of galaxy B. This suggests a radio-quiet AGN (e.g: 2009ApJ...694..235P). Radio-quiet AGN are often associated to spiral galaxies, and share many characteristics with star forming galaxies (2013MNRAS.436.3759B).
@@
Given the evidence of the presence of an AGN at the center of galaxy B, it is possible that AT2022rze is associated to changes on the environment or accretion rate of said AGN. Considering the transition from the AGN (galaxy B) to the LINER region (earlier spectra of AT2022rze, see Fig. 11), AT2022rze could be considered a Changing-look AGN (CLAGN). CLAGN is a fairly recent and slowly growing class of transients, but the number of events is still low. Their characteristics are very diverse, and their searches can be considered as an anomaly detection task (e.g., 2021AJ....162..206S). Therefore, systematic analysis of their time-scales and overall light curve and spectral characteristics is scarce. 2023MNRAS.524..188L developed a method to select CLAGN using multi-wavelength observations. Unfortunately, we do not have multi-wavelength coverage of AT2022rze itself. Still, the host X-ray flux measured before the transient went off, is consistent with the pre-activity X-ray fluxes that 2023MNRAS.524..188L measure for their sample of CLAGNs. In addition, the evolution of W1W2 color at the position of AT2022rze that goes from < 0.5 mag before it is detected (see Table 1) to 1.1 mag (see last column of Table 4) for the last observed epoch, is in agreement with the interpretation presented by 2023MNRAS.524..188L, who argue that this effect is produced by a stronger contribution of hot dust produced by the increase in brightness of the AGN.
@@
We argue in Sect. 3.3 that the steep Balmer decrement values are associated to significant host extinction. However, there is evidence that the Balmer decrement can depend on the accretion-rate for some AGNs and CLAGNs, in such cases the steep Balmer decrement could be related to ionizing photons with low flux (2023ApJ...950..106W). The fact that the Balmer decrement evolves with time (see Fig. 12), suggests that other mechanisms are in place in addition to extinction. Still, the fact that coronal lines can be seen during the whole spectra evolution, suggest the presence of high energies being in place.
@@ 5 Conclusions
We have studied AT2022rze to try to determine its nature. The host of AT2022rze seems to be formed of two galaxies (A and B) undergoing a merging event. The transient is located in the center of galaxy B, located South-East of the plane-of-sky geometrical center of the system. AT2022rze exhibits minimal spectral evolution, consistently displaying narrow Balmer lines alongside various high-excitation features, indicative of a persistent high-energy source. The X-ray luminosity observed near the position of the transient before it went off is much lower than that observed in AGNs (e.g: 1978MNRAS.183..129E; 1990ApJ...361..459E; 1996ASPC..103...70S), although we note that the observations are off-set from the position of the transient. Coronal lines present in the spectra of AT2022rze suggest the presence of soft X-ray and UV emission. As coronal lines are found in AGNs, TDEs, and SNe, their presence alone do not serve as a diagnostic. The overall observed spectroscopic and light curve features of the transient are reminiscent of SLSNe II, TDEs and AGNs.
@@
It is worth noting that the sustained presence of coronal lines in the spectra of some transients has led to the proposal of a new class, referred to as coronal line emitters (ECLs). Although this class groups together events that appear to exhibit similar phenomenology, it is not defined by a shared underlying physical mechanism, as the SLSN, TDE, or AGN classes are. Although ECLs have been mostly associated with TDEs (2024MNRAS.528.4775H), it is still not clear what is their association to CLAGNs (2024MNRAS.528.7076C; 2025arXiv250204080C).
We find possible evidence of Bowen resonance-fluorescence (see Sect. 3.2). The number of events displaying Bowen fluorescence features has been steadily increasing and a new class of transients dubbed Bowen fluorescence flares has been proposed (BFFs 2019NatAs...3..242T). Again, this class is defined by observational characteristics rather than by an underlying physical mechanism, and the nature of these transients is still debated (e.g., 2025ApJ...989..173S).
AT2022rze is a luminous ambiguous transient in the nucleus of a merging galaxy. Many such ambiguous events have recently been found and have been dubbed ambiguous nuclear transients (ANTs, 2025MNRAS.537.2024W), which is another phenomenological class. 2025MNRAS.537.2024W propose that ANTs could be TDEs produced by different mechanisms than that of standard TDEs.
@@ Acknowledgements
P.J.P thanks Stephen Thorp, Laureano Martinez, and Alberto Saldana-Lopez for useful discussion. We thank the referee for their constructive comments. Funded by the European Union (ERC, project number 101042299, TransPIre). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. C.L. is supported by DoE award #DE-SC0025599. CPG acknowledges financial support from the Secretary of Universities and Research (Government of Catalonia) and by the Horizon 2020 Research and Innovation Programme of the European Union under the Marie Skłodowska-Curie and the Beatriu de Pinós 2021 BP 00168 programme, from the Spanish Ministerio de Ciencia e Innovación (MCIN) and the Agencia Estatal de Investigación (AEI) 10.13039/501100011033 under the PID2023-151307NB-I00 SNNEXT project, from Centro Superior de Investigaciones Científicas (CSIC) under the PIE project 20215AT016 and the program Unidad de Excelencia María de Maeztu CEX2020-001058-M, and from the Departament de Recerca i Universitats de la Generalitat de Catalunya through the 2021-SGR-01270 grant. SM acknowledges support from the Research Council of Finland project 350458. FP acknowledges support from the Spanish Ministerio de Ciencia, Innovación y Universidades (MICINN) under grant numbers PID2022-141915NB-C21. AAM and CL are partially supported by DoE award # DE-SC0025599 and Cottrell Scholar Award # CS-CSA-2025-059 from Research Corporation for Science Advancement. This work is based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grants No. AST-1440341 and AST-2034437 and a collaboration including current partners Caltech, IPAC, the Weizmann Institute of Science, the Oskar Klein Center at Stockholm University, the University of Maryland, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, IN2P3, University of Warwick, Ruhr University Bochum, Northwestern University and former partners the University of Washington, Los Alamos National Laboratories, and Lawrence Berkeley National Laboratories. Operations are conducted by COO, IPAC, and UW. The ZTF forced-photometry service was funded under the Heising-Simons Foundation grant No. 12540303 (PI: Graham). SED Machine is based upon work supported by the National Science Foundation under Grant No. 1106171. The Gordon and Betty Moore Foundation, through both the Data-Driven Investigator Program and a dedicated grant, provided critical funding for SkyPortal. W. M. Keck Observatory access was supported by Northwestern University and the Center for Interdisciplinary Exploration and Research in Astrophysics (CIERA). This work has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. The Asteroid Terrestrial-impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile. We acknowledge ESA Gaia, DPAC and the Photometric Science Alerts Team (http://gsaweb.ast.cam.ac.uk/alerts). This research has made use of the NASA/IPAC Extragalactic Database, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research has made use of the SVO Filter Profile Service "Carlos Rodrigo", funded by MCIN/AEI/10.13039/501100011033/ through grant PID2020-112949GB-I00 This research has made use of the SVO Filter Profile Service “Carlos Rodrigo”, funded by MCIN/AEI/10.13039/501100011033/ through grant PID2020-112949GB-I00 The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This research has made use of the CIRADA cutout service at cutouts.cirada.ca, operated by the Canadian Initiative for Radio Astronomy Data Analysis (CIRADA). CIRADA is funded by a grant from the Canada Foundation for Innovation 2017 Innovation Fund (Project 35999), as well as by the Provinces of Ontario, British Columbia, Alberta, Manitoba and Quebec, in collaboration with the National Research Council of Canada, the US National Radio Astronomy Observatory and Australia’s Commonwealth Scientific and Industrial Research Organisation. The Legacy Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS; Proposal ID 2014B-0404; PIs: David Schlegel and Arjun Dey), the Beijing-Arizona Sky Survey (BASS; NOAO Prop. ID 2015A-0801; PIs: Zhou Xu and Xiaohui Fan), and the Mayall z-band Legacy Survey (MzLS; Prop. ID 2016A-0453; PI: Arjun Dey). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory (LBNL). The Legacy Surveys project is honored to be permitted to conduct astronomical research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation.